Thomson Trucking has $15 billion in assets, and its tax rate is 30%. Its basic earning power (BEP) ratio is 10%, and its return on assets (ROA) is 6%. What is its times-interest-earned (TIE) ratio? Round your answer to two decimal places.
Answer #1

BEP ratio=EBIT/Total assets

EBIT=(15*10%)=$1.5 billion

ROA=Net income/Total assets

Net income=(15*6%)=$0.9 billion

(EBIT-interest expense)(1-tax rate)=Net income

(1.5-interest expense)(1-0.3)=0.9

1.5-interest expense=0.9/0.7

interest expense=1.5-(0.9/0.7)

=$0.214285714 billion

TIE=EBIT/interest expense

=1.5/0.214285714

=7